Let f(x) = x + 7 and g(x) = x − 4. Find f(x) ⋅ g(x)
Answer by greenestamps(13198)



In this notation, the "dot" simply means to multiply the two expressions.

f(x).g(x) = f(x)*g(x) = (x+7)(x-4) = x^2+3x-28